Water is filled in a rectangular tank of size 3 m × 2 m × 1 m. (a) Find he total force exerted by the water on the bottom surface of the tank, (b) Consider a vertical side of area 2 m × 1 m. Take a horizontal strip of width δx metre in this side, situated at a depth of x metre on this strip. (c) Find the torque of the force calculated in part (b) about the bottom edge of this side. (d) Find the total force by the water on this side. (e) Find the total torque by the water on the side about the bottom edge. Neglect the atmospheric pressure and take g = 10 m s-2.

Grade:upto college level

1 Answers

Kevin Nash
9 years ago
Sol. a) F = A h ρg. b) The force exerted by water on the strip of width dx as shown, dF = p × A = (xρg) × A c) Inside the liquid force act in every direction due to adhesion. di = F × r